OpenStudy (vera_ewing):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). A. -40x + 24 B. -40x + 31 C. -40x + 64 D. -40x + 70

OpenStudy (vera_ewing):

@Loser66 not quite sure how to find f(g(x)). can you help?

OpenStudy (loser66):

\(f(\color{red}{x}) = -4\color{red}{x} +7\) so that , to find \(f(\color{blue}{g(x)})\) You just replace \(\color{red}{x}\) by \(\color{ blue}{g(x)}\)

OpenStudy (loser66):

show your try

OpenStudy (vera_ewing):

f(10x-6)=-4x+7 10xf-6f=-4x+7

OpenStudy (loser66):

why do you just replace x by g(x) on the left hand side? How about the right hand side?

OpenStudy (vera_ewing):

Oh okay sorry. Hold on.

OpenStudy (vera_ewing):

I got 10xf-6f=-40x+31

OpenStudy (vera_ewing):

@Loser66 so I guess that answer would be B, right?

OpenStudy (loser66):

yup but the left hand side, just let it there. I mean f(g(x)) = -40x +31 Don't replace g(x) there. YOu are not wrong, just answer directly to the question : "What is f(g(x) " got me?

OpenStudy (vera_ewing):

No, I don't understand...

OpenStudy (loser66):

f(g(x)) = f(10x -6) = -4(10x-6) +7 = -40x +31 YOur calculating is correct. Just NOTATION, To make it clearer, for example, if you are a son of Michael and your name is Vera. Instead of call you " The son of Michael" , people call you Vera. Exactly this situation, instead of f ( 10x-6) and you have the name of 10x -6 is g(x), so you just call it f (g(x)) , and calculate it as above

OpenStudy (vera_ewing):

Ohh ok thanks! I get it now :)

OpenStudy (loser66):

:)